OpenStudy (anonymous):

Snowy's Snow Cones has a special bubble gum snow cone on sale. The cone is a regular snow cone that has a spherical piece of bubble gum nested at the bottom of the cone. The radius of the snow cone is 2 inches, and the height of the cone is 3 inches. If the diameter of the bubble gum is 0.5 inches, what is the closest approximation of the volume of the cone that can be filled with flavored ice? Use 3.14 for pi. 22.97 cubic inches 12.50 cubic inches 13.24 cubic inches 12.60 cubic inches

OpenStudy (anonymous):

I know the volume is 12.6 but i don't know how to do the rest of this

OpenStudy (anonymous):

@emcrazy14

OpenStudy (anonymous):

@AravindG @agent0smith @arabpride @aprilann143 @AuroraB @bek_love @baby456 @B.Sanders @Bsho1997 @CrazyCountryGirl @calculusfunctions @ChipperJay* @dinmosiaren @Darkesage

OpenStudy (anonymous):

\[\left( \pi2^2*\frac{ h }{ 3 } \right)-\left( \frac{ 4 }{ 3 }\pi.25^3 \right)\]= V

OpenStudy (anonymous):

h=3

OpenStudy (anonymous):

im so lost

OpenStudy (anonymous):

you have to find the volume of the piece of gum and subtract it from the volume of the cone.

OpenStudy (anonymous):

ok for the volume of the gum i got 33.5

OpenStudy (anonymous):

um did you use the equtaion i gave you?

OpenStudy (anonymous):

you got way to big of a number

OpenStudy (anonymous):

nvm i got 0.065

OpenStudy (anonymous):

thats a lot better

OpenStudy (anonymous):

then i got 12.57

OpenStudy (anonymous):

so what do i do after that

OpenStudy (anonymous):

divide 12.57 by .065?

OpenStudy (anonymous):

no subtract

OpenStudy (anonymous):

why would you divide?

OpenStudy (anonymous):

I got it nvm lol thank you for your help and patience

OpenStudy (anonymous):

your welcome :)